The equivalent capacitance of two capacitors in parallel is greater than the capacitance of either one alone.
Answer #1

Yes , The equivalent capacitance of two capacitors in parallel is greater than the capacitance of either one alone because values of capacitors are added in parallel so definitely it will be greater than either one alone

example :-

Two capacitors are connected in parallel hence the capacitors values are C1 = 2\muF , C2 = 3 \mu F

Equivalent capacitance Ceq = C1 + C2

Ceq = 2+3 = 5\muF

Note :-

Here it was proved that Ceq (5\mu​​​​​​F ) > ( either of C1 or C2 )